[Hallc_running] [New Logentry] RC Update - 07/04/22

sjdkay at jlab.org sjdkay at jlab.org
Mon Jul 4 13:00:02 EDT 2022

Logentry Text:
Some updates -

Accelerator Updates

- Beam has been limited to ~40 uA all weekend, fairly unstable but we've slowly progressed on some settings
- There's a plan to address multiple potential causes of the low current issues on a maintenance day on Wednesday, RF work will be done and a power supply in the injector which may be introducing 60Hz noise is being replaced
- We'll plan for a sweep and a controlled access on Wednesday during the accelerator maintenance, we'll need it to address a few miscellaneous issues in the hall. Speaking of...

Hall Updates

- Garth H noticed a water leak near the HMS hut. This seemed to grow a bit over an hour or so which Brad confirmed with a nice video - https://logbooks.jlab.org/entry/4010658 - I called the tech on call. Rob came in to take a closer look on the cameras - https://logbooks.jlab.org/entry/4010670 - He thinks it's probably an AC tank on the other side of the HMS and it isn't anything to be too concerned about. It will need to be addressed next time we have an access (which will be on Wednesday). It looks to have grown a bit more overnight - https://logbooks.jlab.org/entry/4010893
- Slightly more concerning is an issue we had with the target overnight - https://logbooks.jlab.org/entry/4010720 - Both fans read at 0Hz. Target on call (Dave M) was called and had a look, from what he could see, everything on the target still seems fine so it's probably just a communication issue. However, we will need to get this fixed as soon as there's another access. Nominally, this is Wednesday, *BUT*, I think this seems a little more pressing than the water issue. I would be happy to request an access tomorrow to address this. I'll check with the target experts on this.
- Not critical, but the laptop in the HMS hut needs resetting - https://logbooks.jlab.org/entry/4010673 - Coincidentally, this seems to have happened at the same time as the water leak. Probably not connected unless it was a Gamer(TM) water cooled laptop.
- Another HMS Q1 interlock related to coil 4 this morning - https://logbooks.jlab.org/entry/4010831

Physics Updates

- We finally finished the Q2 = 6.00, W = 2.40, x = 0.55, High Epsilon, SHMS Center LD2+ setting - https://logbooks.jlab.org/entry/4010696
- We then moved to the SHMS left setting and did the dummy, we're finishing up LH2 now. The LD2 left setting is being skipped.
- The current setting will finish around 15:00-16:00, and then we'll move on to the Q2 = 6 Fpi setting. We're starting on SHMS centre (7.65) to avoid the need for a hall access on a holiday.

That's it from me for today, have a great 4th of July!


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4010897
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220704/84980f01/attachment.html>

More information about the Hallc_running mailing list